# Aspects of the chiral crossover transition in (2+1)-flavor QCD with Möbius domain-wall fermions

### v1

@@ Abstract
The non-singlet part of the chiral symmetry in QCD with two light flavors is known to be restored through a crossover transition at a pseudo-critical temperature. However, the temperature dependence of the singlet part of the chiral symmetry and whether it is effectively restored at the same temperature is not well understood. Using (2+1)-flavor QCD configurations generated using the Möbius domain-wall discretization on an $N_{\tau}=8$ lattice, we construct suitable observables where the singlet and non-singlet chiral symmetries are disentangled in order to study their temperature dependence across the crossover transition. From the peak of the disconnected part of the chiral susceptibility, we obtain a pseudo-critical temperature $T_{pc}=158.7{}_{{}\mathop{-}2.3}^{{}\mathop{+}2.6}\,\mathrm{MeV}$ where the non-singlet part of the chiral symmetry is effectively restored. From a calculation of the topological susceptibility and its temperature dependence we find that the singlet $\mathrm{U}_{A}(1)$ part of the chiral symmetry is not effectively restored at $T\lesssim 186\,\mathrm{MeV}$.
@@ I Introduction
Two important physical phenomena, spontaneous chiral symmetry breaking and confinement are responsible for the remarkable features in the phase diagram of strongly interacting matter described by quantum chromodynamics (QCD) [1, 2]. While other more exotic phases are anticipated from the study of model quantum field theories which share the same symmetries with QCD, the transition from a phase of color-singlet hadrons to a phase of quasi-particles carrying color degrees of freedom at small baryon densities [3, 4, 5, 6] is extensively studied using first-principle lattice techniques. It is now established that these two phases of QCD are related by a smooth crossover [7, 8, 9]. In the limiting case of massless quarks, chiral symmetry is restored through a true phase transition, and the corresponding critical temperature $T_{c}$ has been estimated [10, 11]. However, its universality class has not been established yet.
@@
For two massless flavors of quarks, QCD exhibits a $\mathrm{U}_{L}(2)\times\mathrm{U}_{R}(2)$ chiral symmetry. Its non-singlet subgroup $\mathrm{SU}_{L}(2)\times\mathrm{SU}_{R}(2)$ is spontaneously broken to $\mathrm{SU}_{V}(2)$ in the hadronic phase, giving rise to three pseudo-Goldstone modes: the pions, which are much lighter than the nucleons. The singlet $\mathrm{U}_{A}(1)$ subgroup of the chiral symmetry group does not correspond to an exact symmetry in QCD and is broken due to quantum effects arising from gauge interactions [12, 13, 14]. However, this anomalously broken symmetry is still believed to affect the nature of the chiral phase transition of QCD at zero baryon density with two light quark flavors [15]. From renormalization group studies of model quantum field theories with the same symmetries as QCD, two interesting scenarios arise: (1) If the $\mathrm{U}_{A}(1)$ symmetry is approximately restored at $T_{c}$, then the phase transition from the hadronic phase to the quark-gluon plasma phase is expected to be of first order [15, 16] or of second order in the $\mathrm{U}_{L}(2)\times\mathrm{U}_{R}(2)/\mathrm{U}_{V}(2)$ universality class [17, 18]. (2) If the magnitude of the $\mathrm{U}_{A}(1)$ symmetry breaking term is comparable to its zero temperature value at $T_{c}$, then the phase transition is of second order with $\mathrm{O}(4)$ critical exponents [15, 16, 17, 18, 19]. In such model studies the coefficient of the $\mathrm{U}_{A}(1)$ breaking term is a parameter; its magnitude can be estimated only from non-perturbative studies of QCD. It is an ongoing effort to determine the symmetry group near $T_{c}$ which defines the universality class of the chiral phase transition. The lattice regularization is the only practical method for this task and is studied intensively by different groups using different fermion discretizations and methods  [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
QCD thermodynamics on the lattice is typically studied using different variants of either Wilson or staggered fermion discretization schemes, both of which do not possess exact chiral symmetry on the lattice. At finite lattice spacing, the Wilson Dirac operator breaks chiral symmetry explicitly, whereas the staggered fermion discretization reduces the symmetry group to a remnant $\mathrm{U}(1)\times\mathrm{U}(1)$ part of the continuum chiral symmetry group, which introduces additional lattice cut-off effects, the so-called taste violation effects. In both cases, exact chiral and flavor symmetries are only recovered in the continuum limit. It is therefore important to study effects of chiral symmetry in QCD on the lattice using fermion discretizations where the chiral and the continuum limits are disentangled.
@@
To address issues related to the restoration of chiral symmetry in lattice QCD calculations prior to taking a continuum limit, it is most suitable to use fermions with exact chiral symmetry on the lattice. Overlap fermions [35, 36] are the only known lattice fermion discretization which has an exact chiral and flavor symmetry on the lattice. However, practical lattice QCD calculations with overlap fermions at finite temperature in the thermodynamic limit are numerically challenging due to the occurrence of exact zero modes [37, 38]. A closely related approach is the domain-wall fermion discretization [39], which realizes exact chiral symmetry by introducing an extra dimension which is infinitely large and a non-trivial defect-like potential. In lattice simulations where the extent $L_{s}$ of this fifth dimension is finite, a small mixing between the left- and right-handed chiral modes is inevitable and is quantified by the so-called residual mass $m_{\mathrm{res}}$. The violation of chiral symmetry can be parametrized as $\mathrm{e}^{-m_{\mathrm{res}}L_{s}}/L_{s}$  [40] and an exact chiral symmetry is recovered for an infinite fifth dimension, independently from the continuum limit of QCD.
@@
In this work we use a variant of the domain-wall discretization, known as Möbius domain-wall fermions (MDWF) [41, 42], which is optimized to reduce the residual mass even on lattices with a moderately large extent of the fifth dimension. The chiral crossover transition in (2+1)-flavor QCD has been studied earlier using the MDWF discretization for physical light and strange quark masses on an $N_{\tau}=8$ lattice [9]. Several studies have been performed on the fate of chiral symmetry near the crossover transition, using overlap fermions with fixed topology and MDWF in two- as well as (2+1)-flavor QCD, which report an effective restoration of the axial $\mathrm{U}_{A}(1)$ symmetry at the corresponding pseudo-critical temperature [43, 24, 28, 44]. On the other hand, studies for (2+1)-flavor QCD with domain-wall fermions in Ref. [22, 45] and also with MDWF [9] have demonstrated that the $\mathrm{U}_{A}(1)$ symmetry remains strongly broken at this temperature, and more recently it has also been noted that the chiral susceptibility in the crossover region receives a sizeable contribution from the singlet $\mathrm{U}_{A}(1)$ [30]. A recent study in (2+1)-flavor QCD with the HISQ action reports an effective restoration of the $\mathrm{U}_{A}(1)$ symmetry only at $T\gtrsim 180$ MeV after performing a careful continuum limit extrapolation [32]. It is thus imperative to revisit the problem through the use of fermions with exact chiral symmetry on the lattice, as we do in this work.
@@ II Theoretical Background
Since the $\mathrm{U}_{A}(1)$ subgroup of the chiral symmetry group is anomalous, there is no corresponding order parameter. For this reason, we focus on hadronic correlation functions integrated over spacetime, and analyze how these are sensitive to the singlet and non-singlet part of the chiral symmetry. In QCD with two light quark flavors, the four possible meson correlation functions integrated over the spacetime volume are:
@@
where the meson operators $\sigma(x)=\overline{\psi}_{\ell}(x)\,\psi_{\ell}(x)$ and $\delta^{i}(x)=\overline{\psi}_{\ell}(x)\,\tau^{i}\psi_{\ell}(x)$ denote the scalar singlet and iso-triplet states and $\eta(x)=i\overline{\psi}_{\ell}(x)\,\gamma^{5}\,\psi_{\ell}(x)$ and $\pi^{i}(x)=i\overline{\psi}_{\ell}(x)\,\tau^{i}\,\gamma^{5}\,\psi_{\ell}(x)$ denote the pseudo-scalar singlet and iso-triplet states, respectively. Here $\psi_{\ell}$ represents the field operator corresponding to the two degenerate light quark flavors with mass $m_{\ell}$, and $\tau$ represents the isospin operator, which for the $\mathrm{SU}(2)$ flavor group is given by the Pauli matrices.
@@
Furthermore, the chiral Ward identities relate the single-flavor chiral condensate $\langle\bar{\psi}\psi\rangle_{\ell}$ to the iso-triplet pseudo-scalar susceptibility by means of
@@
for a quark mass $m_{f}$, with $s,s^{\prime}=1,2,\cdots,L_{s}$, and where $P_{\pm}=\frac{1\pm\gamma_{5}}{2}$ are the left- and right-chiral projection operators, $D_{W}\equiv D_{W}(-M_{\mathrm{5D}})$ denotes the 4D Wilson Dirac operator with the domain-wall height $M_{\mathrm{5D}}$, and $b_{5}$ and $c_{5}$ are the Möbius parameters, which are tuned to reduce chiral symmetry breaking at fixed $L_{s}$. In the case of the HISQ discretization, the derivatives $\partial_{m_{f}}M_{f}$ in Eqs. (8) and (9) are simply identity matrices with a factor of $1/4$. The restoration of the non-singlet part of chiral symmetry implies that [22, 45, 27]
@@
Here the trace runs over spacetime, spinor and color indices. The topological tunnelings generate a finite topological charge on each gauge configuration and therefore lead to the violation of charge-parity symmetry. However, since QCD is a CP-even theory and physical results are obtained from averages over a large set of gauge configurations, the resulting topological charge, calculated after averaging over the full ensemble of configurations, vanishes and its fluctuations, given by $\chi_{\text{top}}$, are finite. From lattice studies in (2+1)-flavor QCD it is known that the value of the topological susceptibility is consistent with the prediction from chiral perturbation theory [47] below the pseudo-critical temperature and has a non-trivial temperature dependence up to twice its value [27]. At higher temperatures, i.e. beyond $2.5$ to $3$ times the pseudo-critical temperature, its temperature dependence is well described by a dilute instanton gas approximation (DIGA) [27, 48, 49, 5]. Hence, when the non-singlet part of chiral symmetry is restored, the relation
@@
is exactly satisfied. This has been demonstrated within the HISQ discretization scheme in (2+1)-flavor QCD in the continuum limit [27]. Thus, the observable $m_{\ell}^{2}\,\chi_{\text{disc}}-\chi_{\text{top}}$ can be studied as a measure of the breaking of the non-singlet part of chiral symmetry.
For the anomalous $\mathrm{U}_{A}(1)$ symmetry, an effective restoration would imply $\chi_{\pi}=\chi_{\delta}$ [50]. Therefore, the difference between the two integrated correlation functions, $\chi_{\pi}-\chi_{\delta}$, is sensitive to the amount of $\mathrm{U}_{A}(1)$ symmetry breaking. Furthermore, when the non-singlet part of chiral symmetry is restored, then $\chi_{\pi}=\chi_{\sigma}$, and as a result, the observable $\chi_{\pi}-\chi_{\delta}$ is equal to $1/2\,\chi_{\text{disc}}$. In turn, the amount of $\mathrm{U}_{A}(1)$ symmetry breaking in this phase is related to $\chi_{\text{top}}$, because $\chi_{\text{top}}=\chi_{\text{5,disc}}$. Since the temperature dependence of $\chi_{\text{top}}$ can be characterized using DIGA up to asymptotically high temperatures, the $\mathrm{U}_{A}(1)$ symmetry remains broken, reflecting its anomalous nature. A non-trivial temperature dependence of $\chi_{\text{top}}$ which cannot be explained within DIGA will thus denote a non-trivial breaking of the singlet $\mathrm{U}_{A}(1)$ part of chiral symmetry. While the above relations are strictly valid in the continuum for two flavors of quarks in the massless limit, we study how well these observables, which are sensitive to the singlet and non-singlet parts of the chiral symmetry, represent the physics of the chiral phase transition on a finite lattice using MDWF.
@@ III Numerical Methods and Strategy
In this work we have generated gauge configurations for (2+1)-flavor QCD using the MDWF discretization for fermions and the Iwasaki gauge action for the gauge links. To remove localized dislocations in the gauge field, which contribute to the magnitude of the residual mass, we weighted the fermion determinant with the dislocation suppressing determinant ratio (DSDR) [51]. We performed simulations for a fixed lattice size, where the numbers of sites along the spatial and temporal directions were set to $N_{\sigma}=32$ and $N_{\tau}=8$, respectively. The light ($m_{\ell}$) and strange ($m_{s}$) quark mass, the residual mass $m_{\text{res}}$, the extent of the fifth dimension $L_{s}$ and the Möbius parameters $b_{5}$ and $c_{5}$ were fixed following an earlier work using the same setup [9]. The quark masses were fixed such that the pion mass has a value of $129.2(5)\,\mathrm{MeV}$ and the kaon mass is $462.5(5)\,\mathrm{MeV}$, with the ratio of the strange to light quark mass being $m_{s}/m_{\ell}\approx 27$ for all temperatures studied. The earlier study [9] reported a pseudo-critical temperature of $T_{pc}=155(8)(1)\,\mathrm{MeV}$; in this work we perform a more precise determination of it. We have also performed new optimizations within the hybrid Monte-Carlo (HMC) algorithm to achieve a speed-up in the generation of gauge configurations for this study. First, we have used six steps of Hasenbusch mass preconditioning, instead of five as used in the previous work, such that the mass ratios are spaced evenly. This accelerates the inversion of the Dirac matrix for temperatures $T\lesssim T_{pc}$, corresponding to the values $\beta=1.671,1.689,1.707$ of the strong coupling. Furthermore, for $T\gtrsim T_{pc}$, we used four integration steps per HMC trajectory for fermions and $32$ for the gauge links, as opposed to the previously used $48$ integration steps for the latter. With these optimizations, the time required to generate each gauge configuration was reduced by approximately $20\%$, with a typical acceptance rate of about $90\%$ for $T\lesssim T_{pc}$, and of more than $97\%$ for $T>T_{pc}$. The gauge configurations were saved after every fifth HMC step of unit length to reduce the auto-correlation between the generated configurations. Whenever possible, we combined in our analysis the data sets from the earlier analysis [9] and this work. The parameters used for the generation of gauge configurations and the studied temperature range are summarized in Table 1. The temperature values have been fixed using the $r_{0}$ scale, following the earlier work [52]. The numbers of MDWF configurations generated at each temperature are summarized in Table 2. We have estimated the residual mass on these configurations and then combined it with the earlier results [52] for the same quantity, final results of which are also tabulated in Table 2.
@@
In order to estimate the topological charge on each configuration, we have determined the number of zero modes of the overlap Dirac operator, which is used as a probe on the MDWF gauge ensembles. The overlap Dirac operator has an index [53] which is related to the topological features of the underlying gauge configurations if the gauge links are sufficiently smooth [54]. The average value of the topological charge averaged over the available ensembles is consistent with zero, and we have calculated the mean fluctuation of the topological charge to estimate its susceptibility. The overlap Dirac operator was realized such that on each configuration the Ginsparg-Wilson relation and the realization of the sign-function was implemented with a precision of $\lesssim 10^{-9}$. Furthermore, the zero modes were unambiguously identified from their chiralities: a typical zero mode always exhibits a chiral charge of ±$\pm 1$, with a typical precision of $10^{-4}$, whereas near-zero modes come as pairs with equal and opposite chirality. For the estimation of the gauge average, as well as its error, we used the jackknife method, where we have determined the optimal number of jackknife bins for each observable and at each $\beta$-value according to the saturation of the jackknife error estimate.

### v2

@@ Abstract
The nonsinglet part of the chiral symmetry in quantum chromodynamics (QCD) with two light flavors is known to be restored through a crossover transition at a pseudocritical temperature. However, the temperature dependence of the singlet part of the chiral symmetry and whether it is effectively restored at the same temperature is not well understood. Using (2+1)-flavor QCD configurations generated using the Möbius domain-wall discretization on an $N_{\tau}=8$ lattice, we construct suitable observables where the singlet and nonsinglet chiral symmetries are disentangled in order to study their temperature dependence across the crossover transition. From the peak of the disconnected part of the chiral susceptibility, we obtain a pseudocritical temperature $T_{pc}=158.7{}_{{}\mathop{-}2.3}^{{}\mathop{+}2.6}\,\mathrm{MeV}$ where the nonsinglet part of the chiral symmetry is effectively restored. From a calculation of the topological susceptibility and its temperature dependence we find that the singlet $\mathrm{U}_{A}(1)$ part of the chiral symmetry is not effectively restored at $T\lesssim 186\,\mathrm{MeV}$.
@@ I Introduction
Two important physical phenomena, spontaneous chiral symmetry breaking and confinement are responsible for the remarkable features in the phase diagram of strongly interacting matter described by quantum chromodynamics (QCD) [1, 2]. While other more exotic phases are anticipated from the study of model quantum field theories which share the same symmetries with QCD, the transition from a phase of color-singlet hadrons to a phase of quasiparticles carrying color degrees of freedom at small baryon densities [3, 4, 5, 6] is extensively studied using first-principle lattice techniques. It is now established that these two phases of QCD are related by a smooth crossover [7, 8, 9]. In the limiting case of massless quarks, chiral symmetry is restored through a true phase transition, and the corresponding critical temperature $T_{c}$ has been estimated [10, 11]. However, its universality class has not been established yet.
@@
For two massless flavors of quarks, QCD exhibits a $\mathrm{U}_{L}(2)\times\mathrm{U}_{R}(2)$ chiral symmetry. Its nonsinglet subgroup $\mathrm{SU}_{L}(2)\times\mathrm{SU}_{R}(2)$ is spontaneously broken to $\mathrm{SU}_{V}(2)$ in the hadronic phase, giving rise to three pseudo-Goldstone modes: the pions, which are much lighter than the nucleons. The singlet $\mathrm{U}_{A}(1)$ subgroup of the chiral symmetry group does not correspond to an exact symmetry in QCD and is broken due to quantum effects arising from gauge interactions [12, 13, 14]. However, this anomalously broken symmetry is still believed to affect the nature of the chiral phase transition of QCD at zero baryon density with two light quark flavors [15, 16, 17]. From renormalization group studies of model quantum field theories with the same symmetries as QCD, two interesting scenarios arise: (1) If the $\mathrm{U}_{A}(1)$ symmetry is approximately restored at $T_{c}$, then the phase transition from the hadronic phase to the quark-gluon plasma phase is expected to be of first order [15, 18] or of second order in the $\mathrm{U}_{L}(2)\times\mathrm{U}_{R}(2)/\mathrm{U}_{V}(2)$ universality class [19, 20]. (2) If the magnitude of the $\mathrm{U}_{A}(1)$ symmetry breaking term is comparable to its zero temperature value at $T_{c}$, then the phase transition is of second order with $\mathrm{O}(4)$ critical exponents [15, 18, 19, 20, 21]. In such model studies the coefficient of the $\mathrm{U}_{A}(1)$ breaking term is a parameter; its magnitude can be estimated only from nonperturbative studies of QCD. It is an ongoing effort to determine the symmetry group near $T_{c}$ which defines the universality class of the chiral phase transition. The lattice regularization is the only practical method for this task and is studied intensively by different groups using different fermion discretizations and methods  [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
@@
QCD thermodynamics on the lattice is typically studied using different variants of either Wilson or staggered fermion discretization schemes, both of which do not possess exact chiral symmetry on the lattice. At finite lattice spacing, the Wilson Dirac operator breaks chiral symmetry explicitly, whereas the staggered fermion discretization reduces the symmetry group to a remnant $\mathrm{U}(1)\times\mathrm{U}(1)$ part of the continuum chiral symmetry group, which introduces additional lattice cutoff effects, the so-called taste violation effects. In both cases, exact chiral and flavor symmetries are only recovered in the continuum limit. It is therefore important to study effects of chiral symmetry in QCD on the lattice using fermion discretizations where the chiral and the continuum limits are disentangled.
@@
To address issues related to the restoration of chiral symmetry in lattice QCD calculations prior to taking a continuum limit, it is most suitable to use fermions with exact chiral symmetry on the lattice. Overlap fermions [38, 39] are the only known lattice fermion discretization which has an exact chiral and flavor symmetry on the lattice. However, practical lattice QCD calculations with overlap fermions at finite temperature in the thermodynamic limit are numerically challenging due to the occurrence of exact zero modes [40, 41]. A closely related approach is the domain-wall fermion discretization [42], which realizes exact chiral symmetry by introducing an extra dimension which is infinitely large and a nontrivial defectlike potential. In lattice simulations where the extent $L_{s}$ of this fifth dimension is finite, a small mixing between the left- and right-handed chiral modes is inevitable and is quantified by the so-called residual mass $m_{\mathrm{res}}$. The violation of chiral symmetry can be parametrized as $\mathrm{e}^{-m_{\mathrm{res}}L_{s}}/L_{s}$  [43] and an exact chiral symmetry is recovered for an infinite fifth dimension, independently from the continuum limit of QCD.
@@
In this work we use a variant of the domain-wall discretization, known as Möbius domain-wall fermions (MDWF) [44, 45], which is optimized to reduce the residual mass even on lattices with a moderately large extent of the fifth dimension. The chiral crossover transition in (2+1)-flavor QCD has been studied earlier using the MDWF discretization for physical light and strange quark masses on an $N_{\tau}=8$ lattice [9]. Several studies have been performed on the fate of chiral symmetry near the crossover transition, using overlap fermions with fixed topology and MDWF in two- as well as (2+1)-flavor QCD, which report an effective restoration of the axial $\mathrm{U}_{A}(1)$ symmetry at the corresponding pseudocritical temperature [46, 26, 30, 47]. On the other hand, studies for (2+1)-flavor QCD with domain-wall fermions in Refs. [24, 48] and also with MDWF [9] have demonstrated that the $\mathrm{U}_{A}(1)$ symmetry remains strongly broken at this temperature, and more recently it has also been noted that the chiral susceptibility in the crossover region receives a sizeable contribution from the singlet $\mathrm{U}_{A}(1)$ [32]. A recent study in (2+1)-flavor QCD with the HISQ action reports an effective restoration of the $\mathrm{U}_{A}(1)$ symmetry only at $T\gtrsim 180$ MeV after performing a careful continuum limit extrapolation [34]. It is thus imperative to revisit the problem through the use of fermions with exact chiral symmetry on the lattice, as we do in this work.
@@ II Theoretical Background
Since the $\mathrm{U}_{A}(1)$ subgroup of the chiral symmetry group is anomalous, there is no corresponding order parameter. For this reason, we focus on hadronic correlation functions integrated over spacetime, and analyze how these are sensitive to the singlet and nonsinglet part of the chiral symmetry. In QCD with two light quark flavors, the four possible meson correlation functions integrated over the spacetime volume are:
@@
where the meson operators $\sigma(x)=\overline{\psi}_{\ell}(x)\,\psi_{\ell}(x)$ and $\delta^{i}(x)=\overline{\psi}_{\ell}(x)\,\tau^{i}\psi_{\ell}(x)$ denote the scalar singlet and isotriplet states and $\eta(x)=i\overline{\psi}_{\ell}(x)\,\gamma^{5}\,\psi_{\ell}(x)$ and $\pi^{i}(x)=i\overline{\psi}_{\ell}(x)\,\tau^{i}\,\gamma^{5}\,\psi_{\ell}(x)$ denote the pseudo-scalar singlet and isotriplet states, respectively. Here $\psi_{\ell}$ represents the field operator corresponding to the two degenerate light quark flavors with mass $m_{\ell}$, and $\tau$ represents the isospin operator, which for the $\mathrm{SU}(2)$ flavor group is given by the Pauli matrices.
@@
Furthermore, the chiral Ward identities relate the single-flavor chiral condensate $\langle\bar{\psi}\psi\rangle_{\ell}$ to the isotriplet pseudo-scalar susceptibility by means of
@@
for a quark mass $m_{f}$, with $s,s^{\prime}=1,2,\cdots,L_{s}$, and where $P_{\pm}=\frac{1\pm\gamma_{5}}{2}$ are the left- and right-chiral projection operators, $D_{W}\equiv D_{W}(-M_{\mathrm{5D}})$ denotes the 4D Wilson Dirac operator with the domain-wall height $M_{\mathrm{5D}}$, and $b_{5}$ and $c_{5}$ are the Möbius parameters, which are tuned to reduce chiral symmetry breaking at fixed $L_{s}$. In the case of the HISQ discretization, the derivatives $\partial_{m_{f}}M_{f}$ in Eqs. (8) and (9) are simply identity matrices with a factor of $1/4$. The restoration of the nonsinglet part of chiral symmetry implies that [24, 48, 29]
@@
Here the trace runs over spacetime, spinor and color indices. The topological tunnelings generate a finite topological charge on each gauge configuration and therefore lead to the violation of charge-parity symmetry. However, since QCD is a $CP$-even theory and physical results are obtained from averages over a large set of gauge configurations, the resulting topological charge, calculated after averaging over the full ensemble of configurations, vanishes and its fluctuations, given by $\chi_{\text{top}}$, are finite. From lattice studies in (2+1)-flavor QCD it is known that the value of the topological susceptibility is consistent with the prediction from chiral perturbation theory [50] below the pseudocritical temperature and has a nontrivial temperature dependence up to twice its value [29]. At higher temperatures, i.e. beyond $2.5$ to $3$ times the pseudocritical temperature, its temperature dependence is well described by a dilute instanton gas approximation (DIGA) [29, 51, 52, 5]. Hence, when the nonsinglet part of chiral symmetry is restored, the relation
@@
is exactly satisfied. This has been demonstrated within the HISQ discretization scheme in (2+1)-flavor QCD in the continuum limit [29]. Thus, the observable $m_{\ell}^{2}\,\chi_{\text{disc}}-\chi_{\text{top}}$ can be studied as a measure of the breaking of the nonsinglet part of chiral symmetry.
For the anomalous $\mathrm{U}_{A}(1)$ symmetry, an effective restoration would imply $\chi_{\pi}=\chi_{\delta}$ [53]. Therefore, the difference between the two integrated correlation functions, $\chi_{\pi}-\chi_{\delta}$, is sensitive to the amount of $\mathrm{U}_{A}(1)$ symmetry breaking. Furthermore, when the nonsinglet part of chiral symmetry is restored, then $\chi_{\pi}=\chi_{\sigma}$, and as a result, the observable $\chi_{\pi}-\chi_{\delta}$ is equal to $1/2\,\chi_{\text{disc}}$. In turn, the amount of $\mathrm{U}_{A}(1)$ symmetry breaking in this phase is related to $\chi_{\text{top}}$, because $\chi_{\text{top}}=\chi_{\text{5,disc}}$. Since the temperature dependence of $\chi_{\text{top}}$ can be characterized using DIGA up to asymptotically high temperatures, the $\mathrm{U}_{A}(1)$ symmetry remains broken, reflecting its anomalous nature. A nontrivial temperature dependence of $\chi_{\text{top}}$ which cannot be explained within DIGA will thus denote a nontrivial breaking of the singlet $\mathrm{U}_{A}(1)$ part of chiral symmetry. While the above relations are strictly valid in the continuum for two flavors of quarks in the massless limit, we study how well these observables, which are sensitive to the singlet and nonsinglet parts of the chiral symmetry, represent the physics of the chiral phase transition on a finite lattice using MDWF.
@@ III Numerical Methods and Strategy
In this work we have generated gauge configurations for (2+1)-flavor QCD using the MDWF discretization for fermions and the Iwasaki gauge action for the gauge links. To remove localized dislocations in the gauge field, which contribute to the magnitude of the residual mass, we weighted the fermion determinant with the dislocation suppressing determinant ratio (DSDR) [54]. We performed simulations for a fixed lattice size, where the numbers of sites along the spatial and temporal directions were set to $N_{\sigma}=32$ and $N_{\tau}=8$, respectively. The light ($m_{\ell}$) and strange ($m_{s}$) quark mass, the residual mass $m_{\text{res}}$, the extent of the fifth dimension $L_{s}$ and the Möbius parameters $b_{5}$ and $c_{5}$ were fixed following an earlier work using the same setup [9]. The quark masses were fixed such that the pion mass has a value of $129.2(5)\,\mathrm{MeV}$ and the kaon mass is $462.5(5)\,\mathrm{MeV}$, with the ratio of the strange to light quark mass being $m_{s}/m_{\ell}\approx 27$ for all temperatures studied. The earlier study [9] reported a pseudocritical temperature of $T_{pc}=155(8)(1)\,\mathrm{MeV}$; in this work we perform a more precise determination of it. We have also performed new optimizations within the hybrid Monte-Carlo (HMC) algorithm to achieve a speed-up in the generation of gauge configurations for this study. First, we have used six steps of Hasenbusch mass preconditioning, instead of five as used in the previous work, such that the mass ratios are spaced evenly. This accelerates the inversion of the Dirac matrix for temperatures $T\lesssim T_{pc}$, corresponding to the values $\beta=1.671,1.689,1.707$ of the strong coupling. Furthermore, for $T\gtrsim T_{pc}$, we used four integration steps per HMC trajectory for fermions and $32$ for the gauge links, as opposed to the previously used $48$ integration steps for the latter. With these optimizations, the time required to generate each gauge configuration was reduced by approximately $20\%$, with a typical acceptance rate of about $90\%$ for $T\lesssim T_{pc}$, and of more than $97\%$ for $T>T_{pc}$. The gauge configurations were saved after every fifth HMC step of unit length to reduce the auto-correlation between the generated configurations. Whenever possible, we combined in our analysis the datasets from the earlier analysis [9] and this work. The parameters used for the generation of gauge configurations and the studied temperature range are summarized in Table 1. The temperature values have been fixed using the $r_{0}$ scale, following the earlier work [55]. The numbers of MDWF configurations generated at each temperature are summarized in Table 2. We have estimated the residual mass on these configurations and then combined it with the earlier results [55] for the same quantity, final results of which are also tabulated in Table 2.
@@
In order to estimate the topological charge on each configuration, we have determined the number of zero modes of the overlap Dirac operator, which is used as a probe on the MDWF gauge ensembles. The overlap Dirac operator has an index [56] which is related to the topological features of the underlying gauge configurations if the gauge links are sufficiently smooth [57]. The average value of the topological charge averaged over the available ensembles is consistent with zero, and we have calculated the mean fluctuation of the topological charge to estimate its susceptibility. The overlap Dirac operator $D_{ov}$ was realized such that on each configuration the Ginsparg-Wilson relation and the realization of the sign-function was implemented with a precision of $\lesssim 10^{-9}$. Furthermore, the zero modes were unambiguously identified from their chiralities: a typical zero mode always exhibits a chiral charge of ±$\pm 1$, with a precision of at least $10^{-4}$, whereas near-zero modes of $D^{\dagger}_{ov}\,D_{ov}$ come as pairs with equal and opposite chirality distinctly different from identity. The magnitudes of the near-zero modes, calculated within this mixed-action formalism, need to be properly renormalized and there is discussion in the literature on this topic [26]; however, the near-zero modes are not relevant for the study presented here and we will address this issue in a separate work.
For the estimation of the gauge average, as well as its error, we used the jackknife method, where we have determined the optimal number of jackknife bins for each observable and at each $\beta$-value according to the saturation of the jackknife error estimate.